# Marco Randoa1, Martin Jamesb, Alessandro Verria, Lorenzo Rosascoa, Agnese Seminarab1

### v1

@@ Acknowledgements
The data we used to train the agents is a set of $2598$ matrices $\{D_{t}\}_{t=1}^{2598}$. Every matrix $D_{t}\in\mathbb{R}^{1225\times 280}$ contains the odor intensity in every position $(i,j)$ i.e. $(D_{t})_{i,j}$ represents the odor intensity in position $(i,j)$ at time $t$. The source of odor is in position $(20,142)$ and, in order to simplify the training, we considered as terminal states every position in a circle centered in the source position and with radius $10$ called source region. Data are obtained from a direct numerical simulation of the Navier-Stokes equations and the equations of transport of the odor. Environments 1 to 4 are derived from simulations of a channel flow described in ref [33]; Environment 5 corresponds to an additional simulation described below. We preprocess the data to eliminate simulation noise by setting to zero every entry of these matrices when they are smaller than a noise level $n_{\text{lvl}}:=3\times 10^{-6}$. Data information are summarized in Table 1.
Table 1: Data information
Feature Value
Shape $1225\times 280$
Number of time slices $2598$
$n_{\text{thr}}$ $3\times 10^{-6}$
Source location $(20,142)$
@@
In environment 5, the odor is advected by a turbulent open channel flow, with three hemispherical obstacles placed on the ground close to the inlet to generate turbulence. The Navier-Stokes equations (1) and advection-diffusion equation for odor transport (2) are solved using a central second order finite difference scheme. The convective terms are discretized in time using an explicit Adams – Bashforth method; and the viscous and diffusion terms using an implicit Crank-Nicolson method [54]. The code is written in Fortran and is GPU parallelized. The channel is divided into $1024\times 256\times 128$ grid points along streamwise, spanwise and wall-normal directions respectively. The corresponding average spatial resolutions are $\Delta x=5\eta,\Delta y=5\eta,\Delta z=4\eta$, where $\eta$ is the Kolmogorov length scale. Three hemispheres of radius $100\eta$ are placed at a distance of $250\eta$ from the inlet on the ground, equally spaced along the spanwise direction. The channel is forced using a constant pressure gradient. For the velocity field, we impose a no-slip boundary condition at the ground and on the obstacles (${\bm{u}}=0$) and a free-slip boundary on top ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$). The velocity field is periodic along the streamwise and spanwise directions. The bulk Reynolds number is 7800. For the odor field, we impose Dirichlet condition ($c=0$) at the ground, on the obstacles and inlet, no-flux ($\partial_{z}c=0$) on top, and outflow along other directions. Similar to the native environment, we choose the Schmidt number to be 1. The odor source is located downstream of the obstacle and centered at [$640\eta$, $640\eta$, $256\eta$] along streamwise, spanwise and wall-normal directions respectively. The odor source has a Gaussian profile with a standard deviation of $8\eta$.
@@
Environment 6 is similar to environment 5 albeit with a higher bulk Reynolds number of 17500. Here, the channel is divided into $2000\times 500\times 200$ grid points and has an average spatial resolution of $\Delta x=\Delta y=\Delta z=5.5\eta$. The odor source has a Gaussian profile centered at [$825\eta$, $1375\eta$, $550\eta$] with a standard deviation of $3\eta$.
$\rho(\frac{\partial\bm{u}}{\partial t}+\bm{u}\cdot\nabla\bm{u})$ $=-\nabla P+\mu\nabla^{2}\bm{u}+\bm{f};$ (1)
$\nabla\cdot\bm{u}$ $=0.$
$\frac{\partial c}{\partial t}+\mathbf{u}\cdot\nabla c=D\nabla^{2}c+s.$ (2)
@@
Each agent stores the odor concentrations detected in the previous $T$ time steps in a vector $\mathbf{M}$. We introduce an adaptive sensitivity threshold function $s_{\text{thr}}(\cdot)$ defined as
@@
where $e_{i}$ denotes the $i$-th canonical base. As explained in the main text, actions are selected using one of two policies according to the current olfactory state $o_{t}$. More precisely, if the olfactory state $o_{t}$ is not the void state, then the ($\epsilon$-greedy) Q-learning policy is used. Formally, let $Q$ be the Q matrix of the agent and let $o_{t}\neq\emptyset$, then the agent plays the action $a_{t}$ such that
@@
$G(i,j)=$ $\langle e^{-\lambda\tau(i,j)}-\frac{\eta}{1-\gamma}(1-e^{-\lambda\tau(i,j)})\rangle$

### v2

@@ 4 Methods and Materials
The data we used to train the agents (Environment 1) is a set of $2598$ matrices $\{D_{t}\}_{t=1}^{2598}$. Every matrix $D_{t}\in\mathbb{R}^{1225\times 280}$ contains the odor intensity in every position $(i,j)$ i.e. $(D_{t})_{i,j}$ represents the odor intensity in position $(i,j)$ at time $t$. The source of odor is in position $(20,142)$ and, in order to simplify the training, we considered as terminal states every position in a circle centered in the source position and with radius $10$ called source region. Data are obtained from a direct numerical simulation of the Navier-Stokes equations and the equations of transport of the odor. Environments 1 to 4 are derived from Simulation 1, a direct numerical simulation of a channel flow described in Rigolli et al. (2022a) and used to develop a POMDP algorithm in Rigolli et al. (2022b), dataset available from Rigolli et al. (2022c). The simulation represents a boundary layer matching the lowest $\sim$1m of the atmosphere and with horizontal dimensions $\sim$1.9$\times$9.5 m and Reynolds number $Re_{\lambda}\sim 1400$ (on the low side of atmospheric Reynolds typically ranging from 1000 to 10000 Gulitski et al. (2007)). Odor snapshots are extracted at a height .$\sim.5\,m∼ .$ from the ground for Environments 1 and 2 and .$\sim.01\,m∼ .$ for Environments 3 and 4 respectively. We preprocess the data to zero every entry of these matrices when they are smaller than a noise level $n_{\text{lvl}}:=3\times 10^{-6}$ (or 0.13 % relative to concentration at the source). The noise level is increased to 0.22% in Environments 2 and 4. Data information are summarized in Table 2 and 3. Levels of intermittency in Figure 1 show that only a thin core region has intermittency larger than 66%, whereas the most challenging regions at the edge of the plume have intermittency under 33%. For reference, experimental values of 25% to 20% were reported for a surrogate odor in the atmospheric boundary layer, along the centerline at 2 to 15 m from the source Murlis and Jones (1981).
Table 2: Gridworld geometry. From Top: 2D size of the simulation, agents that leave the simulation box continue to receive negative reward and no odor; number of time stamps in the simulation, beyond which simulations are looped; number of actions per time stamp; speed of the agent; noise level below which odor is not detected; location of the source on the grid. See Table 3 for the values of the grid size $\Delta x$ and time stamps at which odor snapshots are saved.
Simulation 1 Simulation 2 Simulation 3
2D simulation grid $1225\times 280$ $1024\times 256$ $2000\times 500$
# time stamps $2598$ $5000$ $5000$
# decisions per time stamp 1 1 1
Speed (grid points / time stamp) 10 10 10
$n_{\text{lvl}}$ $3\times 10^{-6}$ $3\times 10^{-6}$ $10^{-4}$
Source location $(20,142)$ $(128,128)$ $(150,250)$
@@
Environments 5 and 6 correspond to horizontal slices at mid height extracted from two additional simulations we performed to corroborate the results (Simulation 2 and 3). In Simulation 2, the odor is advected by a turbulent open channel flow, with three hemispherical obstacles placed on the ground close to the inlet to generate turbulence. The Navier-Stokes equations (1) and advection-diffusion equation for odor transport (2) are solved using a central second order finite difference scheme. The convective terms are discretized in time using an explicit Adams – Bashforth method; and the viscous and diffusion terms using an implicit Crank-Nicolson method Viola et al. (2020). The code is written in Fortran and is GPU parallelized. The channel is divided into $1024\times 256\times 128$ grid points along streamwise, spanwise and wall-normal directions respectively. The corresponding average spatial resolutions are $\Delta x=5\eta,\Delta y=5\eta,\Delta z=4\eta$, where $\eta$ is the Kolmogorov length scale. Three hemispheres of radius $100\eta$ are placed at a distance of $250\eta$ from the inlet on the ground, equally spaced along the spanwise direction. The channel is forced using a constant pressure gradient. For the velocity field, we impose a no-slip boundary condition at the ground and on the obstacles (${\boldsymbol{u}}=0$) and a free-slip boundary on top ($u_{z}=0,\partial_{z}u_{x}=\partial_{z}u_{y}=0$). The velocity field is periodic along the streamwise and spanwise directions. The bulk Reynolds number is 7800. For the odor field, we impose Dirichlet condition ($c=0$) at the ground, on the obstacles and inlet, no-flux ($\partial_{z}c=0$) on top, and outflow along other directions. Similar to the native environment, we choose the Schmidt number to be 1. The odor source is located downstream of the obstacle and centered at [$640\eta$, $640\eta$, $256\eta$] along streamwise, spanwise and wall-normal directions respectively. The odor source has a Gaussian profile with a standard deviation of $8\eta$.
@@
Simulation 3 is similar to Simulation 2, albeit with a higher bulk Reynolds number of 17500. Here, the channel is divided into $2000\times 500\times 200$ grid points and has an average spatial resolution of $\Delta x=\Delta y=\Delta z=5.5\eta$. The odor source has a Gaussian profile centered at [$825\eta$, $1375\eta$, $550\eta$] with a standard deviation of $3\eta$. For Environments 5 and 6, noise level is 0.01% relative to concentration at the source. See Table for a summary of parameters and how they match the physical dimensions of the domain.
Table 3: Parameters of the simulations. From Left to Right: Simulation ID (1, 2, 3); Length $L$, width $W$, height $H$ of the computational domain; mean horizontal speed $U_{b}=\langle u\rangle$; Kolmogorov length scale $\eta=(\nu^{3}/\epsilon)^{1/4}$ where $\nu$ is the kinematic viscosity and $\epsilon$ is the energy dissipation rate; mean size of gridcell $\Delta x$; Kolmogorov timescale $\tau_{\eta}=\eta^{2}/\nu$; energy dissipation rate $\epsilon=\nu/2\langle(\partial u_{i}/\partial x_{j}+\partial u_{j}/\partial x_{i})^{2}\rangle$; wall unit $y^{+}=\nu/u_{\tau}$ where $u_{\tau}$ is the friction velocity; bulk Reynolds number $Re_{b}=U_{b}(H/2)/\nu$ based on the bulk speed $U_{b}$ and half height; magnitude of velocity fluctuations $u^{\prime}$ relative to the bulk speed; large eddy turnover time $T=H/2u^{\prime}$; $\omega_{\text{save}}$ frequency at which odor snapshots are saved ×$\times\tau_{\eta}$. For each simulation, first row reports results in non dimensional units. Second and third rows provide an idea of how non-dimensional parameters match dimensional parameters in real flows in air and water, assuming the Kolmogorov length is 1.5 mm in air and 0.4 mm in water.
Sim ID $L$ $W$ $H$ $U_{b}$ $\eta$ $\Delta x$ $\tau_{\eta}$ $\epsilon$ $y^{+}$ Reb $\frac{u^{\prime}}{U_{b}}$ $T$ $\omega_{\text{save}}\times\tau_{\eta}$
1 40 8 4 23 0.006 0.025 0.01 39 0.0035 11500 15% $64\tau_{\eta}$ 1
air 9.50 m 1.90 m 0.96 m 36 $\frac{\text{cm}}{\text{s}}$ 0.15 cm 0.6 cm 0.15 s 6.3e-4 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.09 cm
water 2.66 m 0.53 m 0.27 m 8.6 $\frac{\text{cm}}{\text{s}}$ 0.04 cm 0.2 cm 0.18 s 3e-5 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.02 cm
2 20 5 2 14 0.004 0.02 0.005 163 0.0038 7830 15% $95\tau_{\eta}$ 5
air 7.50 m 1.875 m 0.75 m 17.5 $\frac{\text{cm}}{\text{s}}$ 0.15 cm 0.75 cm 0.15 s 6.3e-4 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.142 cm
water 2.00 m 0.50 m 0.20 m 3.9 $\frac{\text{cm}}{\text{s}}$ 0.04 cm 0.2 cm 0.18 s 3e-5 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.038 cm
3 20 5 2 22 0.0018 0.01 0.0025 204 0.0012 17500 13% $141\tau_{\eta}$ 2.5
air 16.7 m 4.18 m 1.67 m 30.6 $\frac{\text{cm}}{\text{s}}$ 0.15 cm 0.83 cm 0.15 s 6.3e-4 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.1 cm
water 4.44 m 1.11 m 0.44 m 6.8 $\frac{\text{cm}}{\text{s}}$ 0.04 cm 0.22 cm 0.18 s 3e-5 $\frac{\text{m}^{2}}{\text{s}^{3}}$ 0.03 cm
$\rho(\frac{\partial\boldsymbol{u}}{\partial t}+\boldsymbol{u}\cdot\nabla\boldsymbol{u})$ $=-\nabla P+\mu\nabla^{2}\boldsymbol{u}+\boldsymbol{f};$ (1)
$\nabla\cdot\boldsymbol{u}$ $=0.$
$\frac{\partial z}{\partial t}+\mathbf{u}\cdot\nabla z=D\nabla^{2}z+s.$ (2)
@@
Each agent stores the odor concentrations detected in the previous $T$ time steps in a vector $\mathbf{M}=(z(s(t-T),t-T),...,z(s(t),t))$. We introduce an adaptive sensitivity threshold function $s_{\text{thr}}(\cdot)$ defined as
@@
where $e_{i}$ denotes the $i$-th canonical base. Actions are steps in any of the 4 directions, labeled relative to the mean flow which assumed fixed and known. The gridworld is infinite, in that agents can leave indefinitely. If they exit the simulation box, they continue to receive zero signal and negative reward -0.001. As explained in the main text, actions are selected using one of two policies according to the current olfactory state $o_{t}$. More precisely, if the olfactory state $o_{t}$ is not the void state, then the ($\epsilon$-greedy) Q-learning policy is used. Formally, let $Q$ be the Q matrix of the agent and let $o_{t}\neq\emptyset$, then the agent plays the action $a_{t}$ such that
@@
Circling recovery.
In this case, the recovery strategy consists of adopting a circling behavior Stupski and van Breugel (2024) when the agent is in a void state. The agent keeps in memory two counters, $t_{\text{void}}$ and $T_{\text{change}}$, as well as an action $a_{\text{void}}\in\mathcal{A}$, initialized to 00, $1$, and $e_{1}$, respectively. The first counter represents the consecutive number of void observations, $a_{\text{void}}$ is the action to play when the void state is reached, and $T_{\text{change}}$ is a time threshold that indicates when to switch the action $a_{\text{void}}$. When the agent reaches a void state, it plays the action $a_{\text{void}}$ and increments the counter $t_{\text{void}}$ by one. If $t_{\text{void}}=T_{\text{change}}$, then the agent resets $t_{\text{void}}$ to zero, increases $T_{\text{change}}$ by one, and updates the action $a_{\text{void}}$ as follows
$a_{\text{void}}\leftarrow\left\{\begin{array}[]{cc}e_{2}&\text{if }a_{\text{void}}=e_{1}\\ -e_{1}&\text{if }a_{\text{void}}=e_{2}\\ -e_{2}&\text{if }a_{\text{void}}=-e_{1}\\ e_{1}&\text{if }a_{\text{void}}=-e_{2}\\ \end{array}\right.$
When the agent receives a non-void observation, it resets $t_{\text{void}}$ to 00, $T_{\text{change}}$ to $1$, and $a_{\text{void}}$ to $e_{1}$.
Cast-Surge recovery.
In this case, the agent plays a cast-surge behavior Baker (1990) when reaching the void state. As in the circling recovery strategy, the agent keeps two counters, $t_{\text{void}}$ and $T_{\text{change}}$, as well as an action $a_{\text{void}}\in\mathcal{A}$, initialized to 00, $1$, and $e_{2}$, respectively. At every step in the void state, the agent plays the action $a_{\text{void}}$ and increments the void counter $t_{\text{void}}$ by one. If $t_{\text{void}}=T_{\text{change}}$, the agent takes an upwind step, doubles the value of $T_{\text{change}}$, updates $a_{\text{void}}$ by setting it to $-a_{\text{void}}$, and resets $t_{\text{void}}$ to zero. When the agent receives a non-void observation, it resets $t_{\text{void}}$ to 00, $T_{\text{change}}$ to $1$, and $a_{\text{void}}$ to $e_{2}$.
@@
$G(i,j)=$ $\langle e^{-\lambda\tau(i,j)}-\frac{\sigma}{1-\gamma}(1-e^{-\lambda\tau(i,j)})\rangle$
@@
4.1 Acknowledgements
This research was supported by grants to AS from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No 101002724 RIDING), the Air Force Office of Scientific Research under award number FA8655-20-1-7028, and the National Institutes of Health (NIH) under award number R01DC018789. L. R. and M.R. acknowledge the financial support of the European Research Council (grant SLING 819789), the European Commission (Horizon Europe grant ELIAS 101120237), the US Air Force Office of Scientific Research (FA8655-22-1-7034), the Ministry of Education, University and Research (FARE grant ML4IP R205T7J2KP; grant BAC FAIR PE00000013 funded by the EU - NGEU) and the Center for Brains, Minds and Machines (CBMM), funded by NSF STC award CCF-1231216. M.R. is a member of the Gruppo Nazionale per l’Analisi Matematica, la Probabilità e le loro Applicazioni (GNAMPA) of the Istituto Nazionale di Alta Matematica (INdAM). This work represents only the view of the authors. The European Commission and the other organizations are not responsible for any use that may be made of the information it contains. We thank Francesco Viola for support and discussions regarding computational fluid dynamics as well as Antonio Celani, Venkatesh Murthy, Yujia Qi, Francesco Boccardo, Luca Gagliardi, Francesco Marcolli and Arnaud Ruymaekers for comments on the manuscript.
Data Availability.
The datasets and code used to perform the experiments are available at the following links